Sally is 5 years older then Molly. The product of their age is 50. How old is Molly? Help!
Answer by Cromlix(4381)

Hi there,
Sally is 5 years older then Molly.
Make Molly = x
Sally = x + 5
The product of their age is 50
x(x + 5) = 50
x^2 + 5x - 50 = 0
Factorize
(x - 5)(x + 10) = 0
x + 10 = 0
x = -10 (Disregard as -ve)
x - 5 = 0
x = 5
Molly is 5 years old.
Sally is 10 years old.
Hope this helps:-)